### v2

# Algorithms for Optimizing Acyclic Queries

Abstract
Most research on query optimization has centered on binary join algorithms like hash join and sort-merge join. However, recent years have seen growing interest in theoretically optimal algorithms, notably Yannakakis’ algorithm. These algorithms rely on join trees, which differ from the operator trees for binary joins and require new optimization techniques. We propose three approaches to constructing join trees for acyclic queries. First, we give an algorithm to enumerate all join trees of an $\alpha$-acyclic query by edits with amortized constant delay, which forms the basis of a cost-based optimizer for acyclic joins. Second, we show that the Maximum Cardinality Search algorithm by Tarjan and Yannakakis constructs a unique shallowest join tree, rooted at any relation, for a Berge-acyclic query; this tree enables parallel execution of large join queries. Finally, we prove that any connected left-deep linear plan for a $\gamma$-acyclic query can be converted into a join tree by a simple algorithm, allowing reuse of optimization infrastructure developed for binary joins.
1 Introduction
The query optimizer sits at the heart of a database system. It takes a query as input and generates a plan for efficient execution, allowing users to program declaratively without worrying about evaluation. Among the many relational algebra operators, join has received significant attention in optimization research. Its compositional nature allows for combining information from multiple relations, constructing complex queries from simple ones, and producing an output asymptotically larger than the inputs. The primary challenge is the join ordering problem to find the best arrangement of many join operations. Most existing research has focused on binary join algorithms such as hash join and sort-merge join, but these can produce unnecessarily large intermediates. Recent work has revived interest in optimal join algorithms, notably Yannakakis’ instance-optimal algorithm [44] for acyclic queries, which runs in linear time in the input and output size, $\mathcal{O}(|\textsf{IN}|+|\textsf{OUT}|)$. Its execution is guided by join trees whose nodes are relations, different from traditional binary join plans with relations at the leaves and join operators at the internal nodes. Although the algorithm is optimal regardless of the join tree, the choice of plan can affect practical performance. In this paper, we study the optimization problem in the context of Yannakakis-style algorithms.
A query optimizer typically has two parts: a plan generator and a cost model to assess each plan. This paper focuses on plan generation and presents three approaches:
  • An algorithm that enumerates all join trees of an acyclic query by edits, with amortized constant time delay.
  • A proof that Maximum Cardinality Search [40] yields a unique shallowest join tree for any Berge-acyclic query, enabling parallel execution for very large queries.
  • A characterization of when the algorithm by Hu et al. [20] can convert a binary join plan into a valid join tree, allowing reuse of existing optimizers.
The rest of the paper is organized as follows: Section˜2 discusses related work; Section˜3 introduces relevant concepts and notations; Section˜4 presents the join tree enumeration algorithm; Section˜5 introduces canonical join trees and their construction; Section˜6 discusses the conversion from binary plans to join trees; finally, Section˜7 points to avenues for future work and concludes the paper. All missing proofs can be found in the appendix, which also contains additional figures, tables and algorithms.
2 Related Work
Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [28, 34, 30], cost-based pruning from the top down [11, 15], greedy heuristics [5, 14, 38], and randomized search [36]. Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [21, 23], while others avoid Cartesian products [28, 29]. Our algorithms restrict the query plans to those running in linear time for acyclic queries. In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^1] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other. We later define this as the line graph of the query hypergraph in Section 3. our algorithms find maximum spanning trees of the weighted join graph. Several algorithms for ordering binary joins are based on dynamic programming and tabulate shared structures among different plans [28, 29]. This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner. Our enumeration algorithm avoids redundant work by enumerating join trees by edits, i.e., it outputs the difference between consecutive join trees instead of the trees themselves. Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
Several recent papers have proposed practical implementations of Yannakakis’ algorithm for acyclic queries [2, 20, 42, 47]. For example, Zhao et al. [47] find that different query plans perform similarly, thanks to the optimality of Yannakakis’ algorithm. They adopt a simple heuristic to construct the join tree by picking the largest input relation as the root, and then greedily attaching the remaining relations into the tree. Inspired by this algorithm, we prove in Section˜5 that for Berge-acyclic queries there is a unique shallowest join tree for any given root where the depth of each node is minimized. Furthermore, this tree can be constructed in linear time by Tarjan and Yannakakis’ Maximum Cardinality Search algorithm [40]. Shallow trees are desirable for parallel execution, where the depth of the tree determines the number of sequential steps. Other practical implementations of Yannakakis’ algorithm leverage existing optimizers for binary joins and convert a binary plan into a join tree [2, 20]. In particular, Hu et al. [20] find that every left-deep linear plan encountered in practice can be converted into a join tree by a simple algorithm. This is not surprising, as we will prove in Section˜6 that every connected left-deep linear plan of a $\gamma$-acyclic query must traverse some join tree from root to leaves.
On the theoretical side, attention has been focused on finding (hyper-)tree decompositions to improve the asymptotic complexity of query processing [37, 19, 16]. The general goal is to find a decomposition with small width which can be used to guide the execution of join algorithms. Most algorithms find a single decomposition with minimum width to achieve the optimal asymptotic complexity [37, 19, 16]. Nevertheless, different decompositions with the same width may still lead to different performance in practice, and cost-based optimization remains crucial. For this, Carmeli et al. [6] propose an algorithm to enumerate tree decompositions with polynomial delay. In this paper, we focus on acyclic queries and their join trees, which are precisely the decompositions with width 1. Our enumeration algorithm can generate all join trees by edits with amortized constant delay.
3 Preliminaries
We focus on full conjunctive queries [1] in this paper and identify each query with its hypergraph, where there is a vertex for each variable and a hyperedge for each relation.
Definition 1.
A hypergraph is a tuple $(X,R,\chi)$ where $X$ is a set of vertices, $R$ a set of hyperedges, and $\chi:R\to 2^{X}$ is a function that assigns to each hyperedge a set of vertices.
Without loss of generality, we limit the scope of our discussion to queries which admit connected hypergraphs. For queries with disconnected hypergraphs, we treat each component separately. We do not consider isolated vertices, empty hyperedges or duplicated hyperedges containing the same set of vertices. Therefore, each hyperedge can be identified with its vertex set. We will use $x\in r$ interchangeably with $x\in\chi(r)$ and apply the common set operations to hyperedges. The size of a hypergraph is the sum of all hyperedges $|H|=\sum_{r\in R}|r|$. We let $H|_{x}$ denote the neighborhood of $x$ in $H$, consisting of all hyperedges containing vertex $x$.
Input: Hypergraph $H$, root $r$
Output: Join tree $T_{r}$ rooted at $r$
1
[…]
4while $R\neq\emptyset$ do
5 $i\leftarrow i+1$
6 $r_{i}\leftarrow r$
7 for $x\in r_{i}\cap X$ do
8    for $r^{\prime}\in R:x\in r^{\prime}$ do
9       $p(r^{\prime})\leftarrow r_{i}$
10
11
12 $X\leftarrow X\setminus r_{i}$
13 $R\leftarrow R\setminus\{r_{i}\}$
14 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
Algorithm 1 MCS from $r\in R(H)$
(a) Hypergraph $H_{6}$
(b) Line graph $L_{6}$
(c) $T^{G}$ by GYO
(d) $T^{M}$ by MCS
Figure 1: A hypergraph $H_{6}$, its line graph $L_{6}$, and two join trees $T^{G}$ and $T^{M}$.
Definition 2.
The line graph $L(H)$ of a hypergraph $H=(X,R,\chi)$ is an undirected simple graph $(R,E)$ whose vertices $R$ are the hyperedges of $H$, and there is an edge $(r_{1},r_{2})\in E$ whenever $r_{1}\cap r_{2}\neq\emptyset$. We extend $\chi$ such that $\chi(e)=\chi(r_{1},r_{2}):=r_{1}\cap r_{2}$ for each edge $e=(r_{1},r_{2})\in E$, and define the edge-weight function $\omega(r_{1},r_{2}):=|r_{1}\cap r_{2}|$.
When there is no ambiguity from the context, we simply write $L(H)$ as $L$. The size of a line graph is the sum of all edge weights $|L|=\sum_{e\in E(L)}\omega(e)$. It reflects the number of join clauses in a SQL query with one clause per variable shared between two relations. We let $L|_{x}$ denote the subgraph of $L(H)$ induced by $H|_{x}$.
Example 3.
Figure˜2(a) shows an example hypergraph of size 14. The line graph $L_{6}=L(H_{6})$ is shown in Figure˜2(b). For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph. The edge weight is given by $\omega(Y,U)=|\{a,d\}|=2$. Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$. Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge. The size of the line graph is $|L_{6}|=19$.
Definition 4.
A join tree $T$ of hypergraph $H$ is a spanning tree of $L(H)$ such that $T|_{x}$ is a connected subtree for each $x\in X(H)$. If a certain vertex is specified as the root, $T$ becomes a rooted join tree where the edges are oriented away from the root.
A join tree $T(L)$ is a subgraph of $L$, so we use $R(T)$ to denote the set of tree nodes and $E(T)$ to denote the set of tree edges. Because $T$ is spanning and acyclic, we always have $R(T)\equiv R(L)$ and $E(T)\subseteq E(L)$. The requirement that $T|_{x\in X(H)}$ is connected is also known as the running intersection property [13]. We write $\mathcal{T}(L(H))$ or $\mathcal{T}(H)$ to denote the set of unrooted join trees. The union join graph $\mathcal{U}(L(H))$ or $\mathcal{U}(H)$ is a spanning subgraph of $L(H)$ where the edge set is the union of all join trees $E(\mathcal{U}(H))=\bigcup_{T\in\mathcal{T}(H)}E(T)$. We use $T_{r}$ to denote a join tree rooted at $r\in R(T)$. When there is no ambiguity, we simply write $T$. The depth of a node $r_{i}$ in the rooted tree $T_{r}$, denoted $0pt(T_{r},r_{i})$, is defined as its distance from the root. Join trees can be constructed by a procedure called GYO reduction.
Definition 5.
A GYO reduction order is a sequence of hyperedges $r_{1},r_{2},\ldots,r_{k}$ such that for each $r_{i<k}$, there is some $r_{p>i}$, called the parent of $r_{i}$, such that $\forall r_{j>i}:r_{i}\cap r_{j}\subseteq r_{p}$.
The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree. It generates the join tree $T^{G}$ as shown in Figure˜2(c). Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic. The detailed definitions are given by Appendix˜A in Appendix˜A.
Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40]. Algorithm˜1 shows a simplified pseudocode of MCS. We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in Figure˜2(d) with Appendix˜A in Appendix˜A.
Computation model.
Throughout the paper we assume the Random Access Machine model of computation, where one can allocate an array of size $n$ in $\mathcal{O}(n)$ time. Constant-time operations include accessing and updating an array element, adding or deleting an element in a linked list, and the common arithmetic operations on integers.
4 Enumerating Join Trees
Our strategy for enumerating join trees is based on the following classic result relating join trees of a hypergraph to maximum spanning trees (MSTs) of its line graph:
Proposition 6 (Maier [27]).
A subgraph of $L(H)$ is a join tree of $H$ if and only if it is a maximum spanning tree of $L(H)$ according to the weight function $\omega$.
A naïve approach is to apply an off-the-shelf algorithm for enumerating the maximum spanning trees of $L(H)$. The best known algorithm for this purpose is due to Eppstein [12], and it works by deriving from the input graph $G$ a so-called equivalent graph $G^{\equiv}$. Every spanning tree of $G^{\equiv}$ corresponds to an MST of $G$ and vice versa. Eppstein proves a lower bound of $\Omega(m+n\log n)$ on constructing the equivalent graph from an arbitrary weighted graph with $m$ edges and $n$ vertices. Then to enumerate all $k$ MSTs of $G$, Eppstein applies existing algorithms to enumerate the spanning trees of $G^{\equiv}$. Since there are optimal spanning tree enumeration algorithms that run in $\mathcal{O}(m+n+k)$ [22, 35], the overall time complexity to enumerate MSTs is $\mathcal{O}(m+n\log n+k)$.
The main result of this section is an algorithm for enumerating join trees leveraging the structure of acyclic hypergraphs and their line graphs. In particular:
  • for any $\alpha$-acyclic hypergraph $H$, we can construct an equivalent graph in $\mathcal{O}(|L|)=\mathcal{O}(m+n)$, where $m$ and $n$ are the numbers of edges and vertices in $L(H)$, thus enumerating the join trees in $\mathcal{O}(|L|+k)$.
  • for any $\gamma$-acyclic hypergraph $H$, we can construct an equivalent graph from the union join graph $\mathcal{U}(H)$ which can be built in $\mathcal{O}(|H|+|\mathcal{U}(H)|)$, lowering the overall time complexity of enumeration to $\mathcal{O}(|H|+k)$.
In the rest of this section, we formally introduce the notion of the equivalent graph $G^{\equiv}$. Thereafter, we outline the main algorithm for enumerating join trees, prove its correctness and analyze its complexity. Finally, we consider the case of $\gamma$-acyclic queries.
4.1 Equivalent Graph
The construction of the equivalent graph $G^{\equiv}$ for a given edge-weighted graph $G$, is based on the following local graph modification, sliding transformation.
Definition 7 (Sliding Transformation [12]).
Let $G$ be an edge-weighted graph with weight function $\omega$. Given edges $e^{*}=(u,v)$ and $e=(v,w)$ where $\omega(e)<\omega(e^{*})$, sliding $e$ along $e^{*}$ moves $e$’s end node $v$ to $u$ such that $e=(u,w)$.
(a) $T_{W}$
(b) 3 Edges Slid
(c) $G^{\equiv}$
(d) $T^{\prime}_{W}$
Figure 2: $T_{W}$ is an MST of the line graph in Figure˜2(b) where the black solid lines stand for the tree edges, the dashed line for an MST edge $(S,U)$, the dotted line for a non-MST edge $(T,Y)$ and gray solid lines for the remaining non-tree edges. $G^{\equiv}$ is the resulting equivalent graph where the thick lines highlight the parallel edges.
We can construct an equivalent graph by repeatedly performing sliding transformations. Sliding may create parallel edges, cycles, and self-loops, so the result is a multigraph.
Definition 8 (Equivalent Graph [12]).
Given weighted graph $G$ with MST $T$ rooted, an equivalent graph $G^{\equiv}$ is obtained from $G$ by exhaustively sliding every edge $e=(v,w)\in E(G)$ along another $e^{*}=(u,v)\in T$ as long as $0pt(T,u)<0pt(T,v)$.
Sliding transformation preserves the identity of each edge $e\in E(G)$ as explained in Section˜4.1. Every spanning tree of $G^{\equiv}$ is an MST of $G$ and vice versa.
Theorem 9 (by Eppstein [12]).
The maximum spanning trees of a weighted graph $G$ are in one-to-one correspondence with the spanning trees of its equivalent graph $G^{\equiv}$. [^2] Minimum spanning trees are considered in the original literature [12].
In the following, we use $p(r)$ to denote the parent of $r$ in a given rooted tree, and $c(r)$ to denote its children. We extend the notation of parent to edges as follows: for any edge $e\in E(T)$, its parent $p(e)\in E(T)$ is another tree edge incident to $e$ and closer to the root $r$. The edges incident to the root do not have a parent and are all siblings. Otherwise, the siblings of $e$ are incident to the same parent, denoted as $s(e)=\left\{e^{\prime}\in E(T)\setminus\{e\}\mid p(e^{\prime})=p(e)\right\}$.
Example 10.
Given the 6-clique line graph $L_{6}$ in Figure˜2(b), we find a rooted MST $T_{W}$ as shown in Figure˜2(a). Among all the tree edges shown as black solid lines, we can only apply the sliding transformation to the edge $(T,P)$ whose weight $\omega(T,P)=|\{a\}|=1$ is lighter than its parent tree edge $(W,T)$ with $\omega(W,T)=|\{a,c\}|=2$. We slide along the tree edge $(W,T)$ to the root so that the edge $(T,P)$ becomes $(W,P)$ as shown by the solid curve in Figure˜2(b). Non-tree edges can slide similarly. We consider two examples, $(S,U)$ illustrated with a dashed line and $(Y,T)$ with a dotted line in Figure˜2(a). All other non-tree edges are shown in light gray. We can slide along tree edge $(U,W)$ to the root so that the edge $(S,U)$ becomes $(S,W)$ as shown by the dashed curve in Figure˜2(b). Both ends of $(Y,T)$ can slide along the tree edges to the root so that the edge becomes a self-loop as shown by the dotted loop in Figure˜2(b). This edge will not appear in any spanning tree of $G^{\equiv}$, and therefore not a part of any MST of $L_{6}$. We refer to such an edge as a non-MST edge, as opposed to an MST edge. By applying sliding transformations to a fixpoint, we obtain an equivalent graph $G^{\equiv}$, where there are two sets of parallel edges highlighted by thick lines in Figure˜2(c). For example, the tree edge $(T,P)$ in $T_{W}$ and non-tree edges $(P,U),(P,T),(P,Y)$ are parallel in $G^{\equiv}$ between $P$ and $W$. We can easily verify that each spanning tree of $G^{\equiv}$ corresponds to an MST of $L_{6}$, such as $T^{\prime}_{W}$ in Figure˜2(d).
The choice of the initial spanning tree can affect the structure of the equivalent graph, but the order of sliding transformations performed has no impact [12].
4.2 Enumerating Join Trees of $\alpha$-Acyclic Hypergraphs
The bottleneck of Eppstein’s algorithm for constructing equivalent graphs of general graphs comes from a subroutine that identifies where each edge will eventually slide to. Because each edge can only slide along a heavier edge, it will eventually be “blocked” by a lighter or equally weighted edge along its path to the root. The subroutine essentially performs binary search to find the blocking edge, leading to the $\log n$ factor in the overall complexity. The key to our improvement is to show that for every acyclic hypergraph $H$, we can construct an “equivalent hypergraph” $H^{*}$ that shares the same structure with $H$ and has the same set of join trees, but $H^{*}$ admits a special join tree with monotonically increasing weight from root to leaf, which enables constant-time identification of the blocking edge.
Definition 11.
A monotonic weight join tree $\hat{T}\in\mathcal{T}(H)$ is a rooted join tree of $H$ such that for any $e\in\hat{T}$ that has a parent $p(e)$, $\omega(e)>\omega(p(e))$.
Example 12.
$H^{*}_{6}$ in Figure˜4(a) is a hypergraph with similar structure to $H_{6}$ in Figure˜2(a). $H^{*}_{6}$ differs from $H_{6}$ by one vertex $d^{\prime}$ and admits the same set of join trees. Therefore, finding an equivalent graph $G^{\equiv}(H^{*}_{6})$ is sufficient for enumerating the join trees of $H_{6}$. $H^{*}_{6}$ also admits a monotonic weight join tree $\hat{T}_{P}$ in Figure˜4(b). This means each tree edge in $\hat{T}_{P}$ is already in place, because it cannot slide along the lighter parent. For each non-tree edge, we can identify its destination in constant time using algorithms for finding lowest common ancestors [3] and level ancestors [4], as detailed in Algorithm˜3.
The example shows we can enumerate join trees of $H$ by constructing $H^{*}$ and its equivalent graph $G^{\equiv}(H^{*})$, then enumerate spanning trees of $G^{\equiv}(H^{*})$. However, constructing $H^{*}$ has a cost that may exceed our target time complexity. We show that, perhaps surprisingly, running Algorithm˜3 directly on $H$ produces the same equivalent graph $G^{\equiv}(H^{*})$, so we can skip the costly construction of $H^{*}$.
Before presenting the algorithm, we first formalize the relationship between $H$ and $H^{*}$ using the concept of hypergraph homomorphisms. Note the definition of homomorphisms is not standardized in the literature, and we choose ours carefully to simplify the proofs. In particular, our definition is equivalent to the homomorphism of incidence graphs [32].
Definition 13 (Hypergraph Homomorphism).
For two hypergraphs $H_{1}$, $H_{2}$, a homomorphism $\mathsf{f}$ is a pair of functions $(\mathsf{f}_{X}:X(H_{1})\rightarrow X(H_{2}),\mathsf{f}_{R}:R(H_{1})\rightarrow R(H_{2}))$ such that:
$\forall r_{1}\in R(H_{1}),x_{1}\in r_{1}:\mathsf{f}_{X}(x_{1})\in\mathsf{f}_{R}(r_{1})$
We write $H_{1}\rightarrow H_{2}$ to denote that there is a homomorphism from $H_{1}$ to $H_{2}$.
When there is no ambiguity from the context, we simply write $\mathsf{f}$ for $\mathsf{f}_{X}$ or $\mathsf{f}_{R}$. Intuitively, homomorphisms preserve relations. A strong homomorphism also reflects relations:
Definition 14 (Strong Homomorphism).
A homomorphism $\mathsf{f}:H_{1}\rightarrow H_{2}$ is called strong if it also satisfies:
$\forall r_{1}\in R(H_{1}),x_{1}^{\prime}\not\in r_{1}:\mathsf{f}_{X}(x_{1}^{\prime})\not\in\mathsf{f}_{R}(r_{1})$
We write $H_{1}\twoheadrightarrow H_{2}$ to denote that there is a strong homomorphism from $H_{1}$ to $H_{2}$.
In this paper we only consider hypergraphs over the same set of edges, [^3] The hypergraphs share the same hyperedge set $R$ but differ in the hyperedge-to-vertices functions $\chi$, i.e., an edge may have different sets of vertices in different hypergraphs. and from now on we assume $\mathsf{f}_{R}$ is the identity function.
We now show that if there is a strong homomorphism $\mathsf{f}:H^{\prime}\twoheadrightarrow H$, then every join tree of $H$ is also a join tree of $H^{\prime}$.
Lemma 15.
If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
We construct $H^{*}$ such that $H^{*}\twoheadrightarrow H$ and $H^{*}$ admits a monotonic weight join tree $\hat{T}$, by first using the MCS algorithm to construct a join tree $T$ of $H$ called the MCS tree. Then, we perturb the weights of $T$ so that they increase monotonically from the root to leaves. Finally, we construct $H^{*}$ such that the perturbed tree is a join tree of $H^{*}$.
The next result shows every MCS tree is already “somewhat monotonic”, in that every edge must contain some variable not in its parent. With abuse of notation we write $e_{1}\subseteq e_{2}$ for $\chi(e_{1})\subseteq\chi(e_{2})$ and use the common set operations directly on edges to reduce clutter.
Lemma 16.
Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
Furthermore, if two edges share any variable not in their parents, they must be siblings.
Lemma 17.
Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
We now show how to derive from an $\alpha$-acyclic hypergraph $H$ an equivalent hypergraph $H^{\prime}$ that preserves $\mathcal{T}(H)$ and admits a monotonic weight join tree $\hat{T}$.
Definition 18 (Vertex Duplication).
Given a hypergraph $H$ and a vertex $x\in X(H)$, duplicating $x$ produces a new hypergraph $H^{\prime}$ that differs from $H$ by only one vertex $\{x^{\prime}\}=X(H^{\prime})\setminus X(H)$ which is added to each $r\in H|_{x}$.
We will refer to such a hypergraph $H^{\prime}$ as a vertex duplication of $H$ or simply a duplicated hypergraph of $H$. We generalize duplication to a set of vertices in the natural way. Importantly, vertex duplication preserves the set of join trees of the original hypergraph.
Lemma 19.
Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
  1. 1.
    $H^{\prime}\twoheadrightarrow H$ and $H^{\prime}\twoheadleftarrow H$
  2. 2.
    $\mathcal{T}(H^{\prime})=\mathcal{T}(H)$
  3. 3.
    $L(H^{\prime})=L(H)$
We are now ready to construct $H^{*}$ by vertex duplication.
Input: MCS tree $T_{r}(H)$
1
2for $r^{\prime}\in c(r)$ do
3 $e\leftarrow(r,r^{\prime});\quad q$.enqueue($e$)
4
5while $q\neq\emptyset$ do
6 $e\leftarrow q.$dequeue()
7 $\Delta\leftarrow\omega(p(e))-\omega(e)+1$
8 if $\Delta>0$ then
9    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
10     make $\Delta$ duplicates of $x$
11
12 for $e^{\prime}\in c(e)$ do
13    $q$.enqueue($e^{\prime}$)
14
15
Algorithm 2 Construction of $H^{*}$
(a) $H^{*}_{6}$
(b) $\hat{T}_{p}$
Figure 3: An equivalent hypergraph $H^{*}_{6}$ and its monotonic weight join tree $\hat{T}_{P}$
Theorem 20.
Given an MCS tree $T$ of an $\alpha$-acyclic hypergraph $H$, there exists a hypergraph $H^{*}$ that
  • admits a monotonic weight join tree $\hat{T}\in\mathcal{T}(H^{*})$ whose edges are in a one-to-one correspondence to those of $T$
  • satisfies $\mathcal{T}(H^{*})=\mathcal{T}(H)$.
Proof.
To prove the existence, we construct such a hypergraph $H^{*}$ by vertex duplication. Section˜4.2 guarantees that $\mathcal{T}(H^{*})=\mathcal{T}(H)$. We only need to show the existence of $\hat{T}\in\mathcal{T}(H^{*})$. Algorithm˜2 constructs $H^{*}$ by vertex duplication in a breadth-first manner from the root of $T$ to make it a monotonic weight join tree in $H^{*}$.
Algorithm˜2 starts with enqueuing each edge incident to the root. Once a tree edge $e\in E(T_{r})$ is dequeued, we check for the number of duplications needed to make it heavier than its parent edge $p(e)$, namely $\Delta=\omega(p(e))-\omega(e)+1$. The duplication is performed only if $\Delta>0$. Section˜4.2 guarantees that $\chi(e)\setminus\chi(p(e))\neq\emptyset$. Therefore, we can always duplicate a vertex $x\in\chi(e)\setminus\chi(p(e))$ for $\Delta$ times such that $\omega(e)>\omega(p(e))$.
Let $e=(r_{u},r_{d})$ where $0pt(T,r_{u})<0pt(T,r_{d})$, $x$ only occurs in the subtree rooted at $r_{u}$. Otherwise, it violates the running intersection property. Therefore, the duplication of $x$ does not affect the weight of any edge closer to the root than $e$ in $T$. By Section˜4.2, duplication in a sibling edge $e^{\prime}\in s(e)$ would not affect the weight of $e$ either, and $\omega(e)>\omega(p(e))$ will continue to hold. The tree has monotonic weights when Algorithm˜2 terminates. ∎
Given a rooted join tree $T$ of a hypergraph $H$, a non-tree edge is $e=(r_{i},r_{j})\in E(L(H))\setminus E(T)$. There is a path in $T$ between $r_{i}$ and $r_{j}$ via their lowest common ancestor $\mathsf{LCA}(r_{i},r_{j})$. We define the LCA edges $\lambda(e)=\lambda(r_{i},r_{j})$ as a set of at most two tree edges on the path and incident to $\mathsf{LCA}(r_{i},r_{j})$. If $r_{i}$, $r_{j}$ are ancestor and child, then $|\lambda(e)|=1$, otherwise, $|\lambda(e)|=2$.
In the process of building $G^{\equiv}(H^{*})$ from a monotonic weight join tree $\hat{T}$ of $H^{*}$, Section˜4.1 and Section˜4.2 guarantee that:
  • no tree edge in $\hat{T}$ needs to be slid;
  • each non-tree edge $e$ can be removed if it is lighter than all $e^{\prime}\in\lambda(e)$;
  • otherwise, the non-tree edge $e$ can be slid directly to be incident to each $e^{\prime}\in\lambda(e)$.
Algorithm˜3 constructs $G^{\equiv}(H^{*})$ from the line graph $L(H)$ weighted by $\omega$ and an MCS tree $T_{r}$. Full details on the constant-time operations on graphs and data structures to support operations performed in Algorithm˜3 can be found in Appendix˜B. As a part of pre-processing, we first conduct a breadth-first search on $\hat{T}_{r}$ to obtain the depth table $0pt$ of each tree node in $\mathcal{O}(|R(\hat{T}_{r})|)$. We will also use the algorithms proposed by Michael Bender [3, 4] to build the data structures in $\mathcal{O}(|R(\hat{T}_{r})|)$ to facilitate the constant-time query of lowest common ancestors (LCAs) and level ancestors (LAs). The constant-time query of depth, LCA and LA allows for finding the $\lambda(e)$ of any non-tree edge $e$ in constant time as shown from Line 3 to Line 3 in Algorithm˜3.
Input: MCS tree $T_{r}(H)$ of line graph $L(H)$ with weight function $\omega$
1 $G^{\equiv}\leftarrow L(H)$
2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
 // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
3 $w^{*}=\omega(e^{*})$
4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
7 if $w^{*}<w_{1}\wedge w^{*}<w_{2}$ then
8      delete $e^{*}$ from $R(G^{\equiv})$
9
10 if $e_{1}=e_{2}\wedge w^{*}=w_{1}$ then
11      slide $r_{j}$ to $r_{1}$
12
13 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}=w_{2}$ then
14      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
15
16 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{1}<w_{2}$ then
17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
18
19 else if $e_{1}\neq e_{2}\wedge w^{*}=w_{2}<w_{1}$ then
20      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
21
22return $G^{\equiv}$
Algorithm 3 buildEGbuildEG
(a) Monotonic Weight Join Tree $\hat{T}_{P}(H^{*}_{6})$
(b) 3 Non-tree Edges Processed
(c) Equivalent Graph $G^{\equiv}(H^{*}_{6})$
Figure 4: Algorithm˜3 on $H^{*}_{6}$ of Figure˜4(a)
Example 21.
We consider $\hat{T}_{P}$ in Figure˜4(b) of an equivalent hypergraph $H^{*}_{6}$ in Figure˜4(a). Three of $\hat{T}_{P}$’s non-tree edges are shown in Figure˜5(a) where the non-MST edges $(Y,W),(Y,T)$ are dotted and the MST edge $(Y,S)$ is dashed. We can identify $\lambda(Y,W)=\{(T,W),(T,U)\}$ in constant time. Because $\omega(Y,W)=|\{a\}|=1$ is lighter than $\omega(T,W)=\omega(T,U)=|\{a,c\}|=2$, it is removed. $(Y,T)$ is also removed, for it is lighter than its LCA edge $(T,U)$. We compare $(Y,S)$ with $\lambda(Y,S)=\{(P,T),(P,S)\}$. Because they have the same weight with $\omega(Y,S)=\omega(P,T)=\omega(P,S)=|\{a\}|=1$, we slide $(Y,S)$ directly to $(T,S)$. The intermediate graph with those three non-tree edges processed is shown in Figure˜5(b). The resulting equivalent graph $G^{\equiv}(H^{*}_{6})$ is shown in Figure˜5(c).
The next result follows immediately from the construction of the equivalent hypergraph.
Lemma 22.
Let $H^{*}=(X^{*},R,\chi^{*})$ be an equivalent hypergraph derived from $H=(X,R,\chi)$ by vertex duplication. For any non-tree edge $e$ of an MCS tree $T$ of $H$,
  • If $\chi(e)=\chi(e^{\prime})$ for some $e^{\prime}\in\lambda(e)$, then $\chi^{*}(e)=\chi^{*}(e^{\prime})$
  • If $\chi(e)\subset\chi(e^{\prime})$ for some $e^{\prime}\in\lambda(e)$, then $\chi^{*}(e)\subset\chi^{*}(e^{\prime})$
Proof.
The duplication of a vertex $x\in\chi(e)$ adds $x^{\prime}$ to both $\chi(e)$ and $\chi(e^{\prime})$. If $\chi(e)=\chi(e^{\prime})$, $\chi^{*}(e)=\chi(e)\cup\{x^{\prime}\}=\chi(e^{\prime})\cup\{x^{\prime}\}=\chi^{*}(e^{\prime})$. A similar argument applies if $\chi(e)\subset\chi(e^{\prime})$. ∎
Section˜4.2 allows us to produce $G^{\equiv}(H^{*})$ by keeping all tree edges of the MCS tree $T(H)$ in place and sliding the non-tree edges directly to their LCA edges. For example, running Algorithm˜1 on the MCS tree $T(H_{6})$ in Figure˜2(d) produces an identical $G^{\equiv}(H^{*}_{6})$ in Figure˜5(c).
Theorem 23.
Let $H$ be an $\alpha$-acyclic hypergraph with weight function $\omega$, line graph $L$ and MCS tree $T$. Let $H^{*}$ be the equivalent hypergraph of $H$ with weight function $\omega^{*}$. Applying Algorithm˜3 to inputs $L$, $T$ and $\omega$ produces the same equivalent graph as applying Algorithm˜3 to inputs $L$, $T$ and $\omega^{*}$.
Proof.
One can readily check that for any non-tree edge $e$ of $T$, the conditions of each if statement in Algorithm˜3 hold under $\omega$ if and only if they hold under $\omega^{*}$ by Section˜4.2. ∎
Theorem 24.
Given line graph $L$ of $\alpha$-acyclic hypergraph $H$ with weight function $\omega$ and an MCS tree $T$ of $H$, Algorithm˜3 returns equivalent graph $G^{\equiv}$ of equivalent hypergraph $H^{*}$ in $\mathcal{O}(m+n)$ where $n$ is the number of vertices and $m$ is the number of edges in $L$.
Proof.
The construction of each data structure $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$ can be done in $\mathcal{O}(m+n)$ [3, 4]. Execution of the rest of the algorithm also takes time $\mathcal{O}(m+n)$, as it requires examining each non-tree edge to either delete it or slide it directly to its LCA edges in the MCS tree. For each non-tree edge, the lookup of at most two LCA edges takes time $\mathcal{O}(1)$, as does the comparison of weights and the direct sliding transformation. ∎
Once we have the equivalent graph $G^{\equiv}$, the existing algorithm allows for enumerating all its spanning trees with an amortized constant delay. The total time complexity of enumerating $|\mathcal{T}(H)|$ given the line graph $L(H)$ as input is $\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
Starting from an $\alpha$-acyclic query $\mathcal{Q}$, Section˜C.1 describes how to convert it into a hypergraph $H$. Section˜C.2 describes how to convert it into a weighted line graph $L(H)$. Both conversions take linear time $\mathcal{O}(|\mathcal{Q}|)$ with respect to the query size as defined in Appendix˜C. Algorithm˜1 constructs an MCS tree $T$ in $\mathcal{O}(|H|)\leq\mathcal{O}(|\mathcal{Q}|)$. With the inputs $L(H)$ and $T$, we can run Algorithm˜3 to construct $G^{\equiv}(H^{*})$ in $\mathcal{O}(m+n)\leq\mathcal{O}(|\mathcal{Q}|)$ (where $n$ is the number of vertices and $m$ is the number of edges in $L(H)$). Finally, we apply the spanning tree enumeration algorithm by Kapoor and Ramesh [22] on $G^{\equiv}(H^{*})$ to enumerate all join trees $|\mathcal{T}(H)|$ in $\mathcal{O}(m+n+|\mathcal{T}(H)|)\leq\mathcal{O}(|\mathcal{Q}|+|\mathcal{T}(H)|)=\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
4.3 $\gamma$-Acyclic Queries
The run time of Algorithm˜3 depends on the size of the line graph which can be quadratically larger than the input hypergraph. But if $H$ is $\gamma$-acyclic, we can bring the total time complexity of enumeration down to $\mathcal{O}(|H|+|\mathcal{T}(H)|)$. The line graph of a $\gamma$-acyclic hypergraph $H$ is its union join graph [25] $L(H)=\mathcal{U}(H)$, i.e., every edge $e\in E(L(H))$ in the line graph is an MST edge. By modifying the MCS algorithm to Algorithm˜7 in Appendix˜F, we can construct a weighted MCS tree $T$ and an unweighted line graph $L(H)$, namely the union join graph, in $\mathcal{O}(|H|+|E(\mathcal{U}(H))|)$. We simplify Algorithm˜3 to Algorithm˜8 in Appendix˜F, which compares the LCA edges $\lambda(e)$ of each non-tree edge $e$ and slides it to construct the equivalent graph $G^{\equiv}(H)$ in $\mathcal{O}(|E(\mathcal{U}(H))|)$. As $\gamma$-acyclicity guarantees that every non-tree edge $e$ is an MST edge, we can use the weights of the LCA edges to determine how to slide $e$. Let $\lambda(e)=\{e_{l},e_{r}\}$. If $e_{l}=e_{r}$, we slide $e$ to be parallel to $e_{l}$. Otherwise, if $e_{l}$ and $e_{r}$ have different weights, we slide $e$ to be parallel to the heavier one. If they have the same weight, we slide $e$ to form a triangle with them.
The number of join trees is lower bounded by the number of non-tree edges in its line graph, namely $|\mathcal{T}(H)|=\Omega(|E(\mathcal{U}(H))|-|E(T)|)$. The total time complexity of enumerating all join trees is $\mathcal{O}(|H|+|E(\mathcal{U}(H))|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
5 The Canonical Join Tree of a Berge-Acyclic Query
An acyclic query can have exponentially many join trees with respect to its size. For example, the line graph of a clique query with $n$ relations is an $n$-clique $K_{n}$ with $n^{n-2}$ join trees by Cayley’s formula [7]. Enumerating all join trees can be prohibitive for large queries. On the other hand, the query optimizer does not need to consider all possible join trees to achieve good performance. For example, the implementation of Yannakakis’ algorithm by Zhao et al. [47] has similar performance on any join tree rooted at the largest relation. An alternative to enumeration is therefore to simply construct an arbitrary join tree for a given root. This can be done in linear time by the Maximum Cardinality Search (MCS) algorithm from a chosen relation as shown in Algorithm˜1. In this section, we show that the MCS algorithm produces a shallowest tree for a Berge-acyclic query, in the sense that the depth of every node in the tree is minimized. We prove that this shallowest tree is unique, and therefore call it the canonical join tree:
Definition 25.
A join tree $T_{r}$ rooted at $r$ is canonical if $0pt(T_{r},r_{i})\leq 0pt(T^{\prime}_{r},r_{i})$ for any other join tree $T^{\prime}_{r}$ rooted at $r$ and any $r_{i}\in R(T_{r})=R(T^{\prime}_{r})$.
A shallow join tree has practical benefits. For example, the depth of the join tree determines the number of sequential steps required in a parallel join algorithm. A shallow join tree tends to be wide and have more leaves, allowing better utilization of indices.
Although Berge-acyclicity was thought to be too restrictive when it was first introduced to database theory [13], we found it to be general enough to cover almost all acyclic queries encountered in the wild. As shown in Table˜1 of Appendix˜F, among 10 45410\,454 queries from five popular benchmarks, $9285$ are $\alpha$-acyclic, and only 8 of these are not Berge-acyclic. In retrospect, this should not be surprising, as most joins in relational databases are over primary/foreign keys. Emerging workloads in graph databases usually involve simple graphs and seldom require composite key joins. A query without composite key joins admits a linear hypergraph, where each pair of hyperedges shares at most one vertex. The following result establishes an equivalence between $\alpha$-acyclicity with linearity and Berge-acyclicity.
Proposition 26.
An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
By Section˜5, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1. Every spanning tree is a maximum spanning tree, therefore a join tree.
Corollary 27.
For a Berge-acyclic hypergraph $H$, any spanning tree of $L(H)$ is a join tree.
In the rest of this section, we prove the existence and uniqueness of the canonical join tree rooted at any relation of a Berge-acyclic hypergraph, and show that it can be constructed by MCS as in Algorithm˜1.
5.1 Existence & Uniqueness of the Canonical Join Tree
The key insight leading to the existence and uniqueness of the canonical join tree is that the line graph of a Berge-acyclic hypergraph is geodetic [31], meaning that there is a unique shortest path between any pair of vertices. To prove this, we first review some relevant concepts and properties of graphs. All graphs below are simple and undirected.
Definition 28.
A cycle in a graph is a sequence of distinct vertices $v_{0},\ldots,v_{k-1\geq 2}$ such that there is an edge between $v_{i}$ and $v_{(i+1)\bmod k}$ for all $1\leq i\leq k$.
A special class of graphs called chordal graphs are intimately related to acyclic hypergraphs:
Definition 29.
A graph is chordal if every cycle of length at least 4 has a chord, i.e., an edge that is not part of the cycle but connects two vertices of the cycle.
The key step in this section is to show the line graph of a Berge-acyclic hypergraph is a block graph, which is a special class of chordal graphs. There are many different characterizations of block graphs [18], and we present a recent one:
Definition 30 (Block Graph [10]).
A graph $G=(R,E)$ with the vertex set $R$ and the edge set $E$ is a block graph if it is chordal and diamond-free.
Removing an edge from a 4-clique $K_{4}$ results in a diamond. Diamond-free requires that no induced subgraph $G|_{R^{\prime}}$ is a diamond, where $R^{\prime}\subseteq R$ is any subset of vertices. We now present the main result relating block graphs to Berge-acyclic hypergraphs:
Lemma 31.
The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], Section˜5.1 implies the following corollary.
Corollary 32.
Let $L$ be the line graph of a Berge-acyclic hypergraph $H$. There is a unique shortest path between any two vertices in $L$.
We are now ready to prove the existence and uniqueness of the canonical join tree.
Theorem 33.
Let $L$ be the line graph of a Berge-acyclic hypergraph $H$, and $\mathsf{P}(r,r^{\prime})$ be the shortest path in $L$ between $r,r^{\prime}\in R(L)$, $T_{r}=\bigcup_{r^{\prime}\in R(L)}\mathsf{P}(r,r^{\prime})$ is the unique canonical join tree for $H$ rooted at $r$.
Proof.
A spanning tree of $L$ is its join tree by Section˜5. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
$T_{r}$ is spanning and connected because it contains a path from $r$ to each $r^{\prime}\in R(L)$.
We prove $T_{r}$ is acyclic by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$. Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$. $R_{0}=\{r\}$ contains only the root. The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic. Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$. Section˜5.1 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$. Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$. Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$. Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a join tree.
The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$. The canonical tree is unique by Section˜5.1. ∎
5.2 Construction of the Canonical Join Tree
Given a Berge-acyclic hypergraph $H$ and $r\in R(H)$ as the root, we can construct the canonical tree $T_{r}(H)$ by running Algorithm˜1. We first present a useful lemma on the structure of the canonical join tree:
Lemma 34.
Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
We now show that the MCS algorithm constructs the canonical join tree.
Theorem 35.
Given a Berge-acyclic hypergraph $H$, Algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
Proof.
Algorithm˜1 labels each hyperedge in $R(H)$ in ascending order from the root $r_{1}=r$ to $r_{n=|R(H)|}$. When Algorithm˜1 has labeled $i\in[n]$ hyperedges, we consider a graph $G_{i}$ whose vertices are the labeled hyperedges $\{r_{1},\ldots,r_{i}\}$ and edges are the parent-child relationships $\{(r_{2},p(r_{2})),\ldots,(r_{i},p(r_{i}))\}$.
To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$. An empty graph is trivially a subtree of any graph. When the root $r$ is labeled as $r_{1}$, this graph of a single vertex is also trivially a subtree of $T_{r}$. Assuming $G_{k}$ is a subtree of $T_{r}$, Algorithm˜1 proceeds to label $r_{k+1}$ and assigns its parent $p(r_{k+1})=r_{p\leq k}$. Suppose for the sake of contradiction that $r_{k+1}$ has a different parent $r_{q}$ in $T_{r}$. Let us consider the possible relationships between $r_{p}$ and $r_{q}$. First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating Theorem˜33. Suppose $r_{q}$ is an ancestor of $r_{p}$. By the running intersection property, $r_{q}\cap r_{k+1}$ must contain $r_{p}\cap r_{k+1}$, but that would imply every vertex in $r_{p}\cap r_{k+1}$ is already marked by $r_{q}$ and the algorithm would not have assigned $r_{p}$ as the parent of $r_{k+1}$. Therefore, $r_{p}$ and $r_{q}$ are not descendants of each other. Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$. Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in Figure˜5 of Appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in Figure˜5. By Section˜5.2, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct. Consider the following possible cases:
Case 1:
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q}$ form a Berge-cycle.
Case 2:
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q}$ form a Berge-cycle.
Case 3:
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
Case 4:
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
All possible cases above contradict Berge-acyclicity, therefore, Algorithm˜1 must correctly assign $r_{p}$ as the parent of $r_{k+1}$. ∎
6 Converting a Binary Join Plan to a Join Tree
A recent approach [20, 2] that converts a binary join plan into a join tree has gained popularity as it allows system builders to leverage existing query optimizers designed for binary join plans. In this section, we focus on an algorithm by Hu et al. [20] that converts left-deep linear join plans into join trees. The algorithm can convert all acyclic queries in standard benchmarks, including those in Table˜1 of Appendix˜F. We prove that the algorithm converts any connected left-deep linear join plan into a join tree if and only if the query is $\gamma$-acyclic. Our result yields a new characterization of $\gamma$-acyclic queries. We formally define binary join plans and describe the algorithm by Hu et al. [20] with Algorithm˜6 in Appendix˜F.
Definition 36.
A left-deep linear plan is a sequence of relations $r_{1},r_{2},\ldots,r_{n}$. It is connected if for each $r_{i\geq 2}$, $\exists r_{j<i}:r_{i}\cap r_{j}\neq\emptyset$.
Query optimizers strive to produce connected plans, to avoid expensive Cartesian products. Many optimizers produce exclusively left-deep linear plans. Plans that are not left-deep are called bushy, and such plans may still be decomposed into left-deep fragments [43].
Given a left-deep linear plan, Hu et al. [20] generate a join tree with Algorithm˜6 in Appendix˜F. The algorithm chooses the first relation $r_{1}$ as the root and iterates through the rest of the plan. For each relation $r_{i\in[2,n]}$, it finds the first relation $r_{j}$ that contains all attributes shared by $r_{i}$ with all previous relations $r_{i}\cap\bigcup r_{k<i}$, and assigns $r_{j}$ as the parent of $r_{i}$. The algorithm constructs a join tree if it finds a parent for each $r_{i}$.
Hu et al. [20] proved that the algorithm succeeds whenever the input plan is the reverse of a GYO-reduction order. They also observed that every left-deep linear plan produced for queries in standard benchmarks is indeed the reverse of a GYO-reduction order. This is not a coincidence, as we show that every connected left-deep linear join plan must be the reverse of a GYO-reduction order if and only if the query is $\gamma$-acyclic.
Theorem 37.
A query is $\gamma$-acyclic if and only if every connected left-deep linear join plan for the query is the reverse of a GYO-reduction order.
The “only if” direction is more difficult. We prove its contrapositive by constructing a $\gamma$-cycle from a query plan that is not the reverse of a GYO-reduction order. We start with a simple observation that directly follows from the definition of GYO-reduction order.
Proposition 38.
If a left-deep linear plan is not the reverse of a GYO-reduction order, then there exists a relation $r_{i}$ that has no parent among $\{r_{j<i}\}$, formally $\neg\exists r_{j<i}:(r_{i}\cap\bigcup r_{k<i})\subseteq r_{j}$. We call $r_{i}$ an orphan, and denote it as $\mathring{r_{i}}$.
We introduce an ordering on relations to compare them in a plan.
Definition 39.
Given a relation $r$, we write $r_{1}\geq_{r}r_{2}$ if $r_{1}\cap r\supseteq r_{2}\cap r$.
We first observe that $\geq_{r}$ is a preorder (reflexive and transitive), but not a partial order (antisymmetry may fail). Moreover, if there is a greatest element $r_{p}$ with respect to $\geq_{r}$, i.e., $\forall r^{\prime}:r_{p}\geq_{r}r^{\prime}$, then $r_{p}$ is a parent of $r$. Those observations lead to the following.
Lemma 40.
Let $\hat{r}$ be a maximal relation with respect to $\geq_{\mathring{r}}$ for an orphan $\mathring{r}$, meaning $\neg\exists r:r>_{\mathring{r}}\hat{r}$, then there is a relation $\bar{r}$ incomparable with $\hat{r}$, i.e., $\hat{r}\not\geq_{\mathring{r}}\bar{r}$ and $\bar{r}\not\geq_{\mathring{r}}\hat{r}$.
Proof.
If every relation is comparable with $\hat{r}$, the maximal relation $\hat{r}$ is the greatest element with respect to $\geq_{\mathring{r}}$, thus a parent of $\mathring{r}$. It contradicts the fact that $\mathring{r}$ is an orphan. ∎
Given a pair of incomparable relations $\hat{r}$ and $\bar{r}$ with respect to $\mathring{r}$, we can find at least one variable in each relation that does not appear in the other relation but appears in $\mathring{r}$, namely $\exists\bar{x}\in\bar{r}\cap\mathring{r}:\bar{x}\notin\hat{r}$ and $\exists\hat{x}\in\hat{r}\cap\mathring{r}:\hat{x}\notin\bar{r}$. We say that $\hat{x}$ is a dangling variable of $\hat{r}$ and $\bar{x}$ is a dangling variable of $\bar{r}$. We are now ready to prove Theorem˜37. Our strategy is to derive a $\gamma$-cycle from any orphan relation, leading to a contradiction.
Proof of Theorem˜37.
To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order. By Section˜6, there is an orphan $\mathring{r}$ in the plan. Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$. Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, Section˜6 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$. There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.
Because the shorter prefix $\mathsf{pre}({t-1})=r_{1},\ldots,r_{t-1}$ is connected, there is a path in $\mathsf{pre}({t-1})$ between $\hat{r}$ and $\bar{r}$. The incomparability guarantees that $\bar{r}\cap\mathring{r}\not\subseteq\hat{r}\cap\mathring{r}$. We can choose a shortest path in $\mathsf{pre}({t-1})$ that connects a variable in $\hat{r}\cap\mathring{r}$ to a variable in $\mathring{r}\setminus\hat{r}$. We now consider two exhaustive cases where each yields a $\gamma$-cycle:
Case 1:
If the shortest path consists of a single relation $\tilde{r}$, $\tilde{r}\not>_{\mathring{r}}\hat{r}$ (because $\hat{r}$ is maximal) and $\tilde{r}\not\leq_{\mathring{r}}\hat{r}$ (because $\tilde{r}$ contains some variable in $\mathring{r}\setminus\hat{r}$). Therefore, $\tilde{r}$ is incomparable with $\hat{r}$. Since $\tilde{r}\cap\hat{r}\neq\emptyset$, $\hat{r},\hat{x},\mathring{r},\tilde{x},\tilde{r},\breve{x}$ form a $\gamma$-cycle where $\hat{x}\in\hat{r}\cap\mathring{r}$, $\tilde{x}\in\tilde{r}\cap\mathring{r}$, and $\breve{x}\in\tilde{r}\cap\hat{r}$.
Case 2:
If the shortest path consists of at least two relations, let $r_{0}=\mathring{r}$, and the path be $r_{1},\ldots,r_{k-1}$. We now show that $r_{0},r_{1},\ldots,r_{k-1}$ form a pure cycle where each $r_{i}$ is only adjacent to $r_{(i-1)\bmod k}$ and $r_{(i+1)\bmod k}$. First, if a $r_{i\in[2,k-2]}$ on the path is adjacent to a $r_{j\in[1,k-1]}$ other than its neighbors $r_{i-1}$ or $r_{i+1}$, the path can be shortened, contradicting the shortest path assumption. Second, if the relation $r_{i\in[2,k-2]}$ is adjacent to $r_{0}$, then $r_{i}$ must contain variables in either $r_{0}\cap\hat{r}$ or $r_{0}\setminus\hat{r}$. We can shrink the path to $r_{1},\ldots,r_{i}$ or $r_{i},r_{k-1}$, which also contradicts the shortest path assumption. Therefore, each $r_{i\in[0,k-1]}$ on the cycle is only adjacent to $r_{(i-1)\bmod k}$ and $r_{(i+1)\bmod k}$. The cycle $r_{0},r_{1},\ldots,r_{k-1}$ is a pure cycle of length $k\geq 3$. A pure cycle is always a $\gamma$-cycle.
This concludes the proof for the “only if” direction.
Now we prove the “if” direction by contrapositive, namely that if a query contains a $\gamma$-cycle, then there exists a connected left-deep linear plan that is not the reverse of a GYO-reduction order. Let the $\gamma$-cycle of length $k\geq 3$ be $r_{0},\ldots,r_{i},\ldots,r_{k-1}$ where $r_{i\in[1,k-2]}$. By definition, there is an $x_{i-1}$ appearing exclusively in $r_{i-1}$ and $r_{i}$ and an $x_{i}$ appearing exclusively in $r_{i}$ and $r_{i+1}$. We modify the original plan by moving $r_{i}$ after $r_{k-1}$ such that $r_{0},\ldots,r_{i-1},r_{i+1},\ldots,r_{k-1},r_{i}$ becomes part of the new query plan.
The prefix $\mathsf{pre}({k-1})=r_{0},\ldots,r_{i-1},r_{i+1},\ldots,r_{k-1}$ is connected. Therefore, the new query plan is connected. However, $r_{i}$ has no parent, because no relation in $\mathsf{pre}({k-1})$ contains both $x_{i-1}$ and $x_{i}$. This implies the plan is not the reverse of a GYO-reduction order. ∎
7 Conclusion and Future Work
We proposed three approaches for constructing join trees. Our enumeration algorithm in Section˜4 generates join trees by edits with amortized constant delay, forming a basis for cost-based optimization of acyclic joins. In Section˜5, we showed that the Maximum Cardinality Search algorithm constructs the unique shallowest join tree for any Berge-acyclic query, supporting very large queries and enabling parallel execution. Finally, in Section˜6, we gave theoretical justification for converting binary join plans to join trees, allowing reuse of existing optimization infrastructure.
Future work includes compact representations of join trees for dynamic programming, as in binary plan optimizers, and the challenging cost estimation for Yannakakis-style algorithms: the random-walk approach [26] models joint probabilities for binary joins, whereas an efficient and accurate solution for semijoins remains to be found.
Our results also raise theoretical questions. Can join tree enumeration achieve worst-case constant delay? We proved Berge-acyclicity sufficient for the existence and uniqueness of the canonical join tree, but it is not necessary, and $\gamma$-acyclicity is insufficient as shown in Figure˜6 of Appendix˜F. What is the precise characterization of hypergraphs that admit a unique canonical join tree for any root, or for some root?
Appendix A Extended Preliminaries
Definition 41 (Hypergraph Acyclicity).
A hypergraph $H$ is:
  • $\alpha$-acyclic if it admits a join tree;
  • $\beta$-acyclic if every subgraph of $H$ is $\alpha$-acyclic;
  • $\gamma$-acyclic if it does not contain any $\gamma$ cycle. A $\gamma$ cycle is a sequence of length $k\geq 3$ of distinct hyperedges and distinct vertices $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that every $x_{i\in[0,k-2]}$ belongs to $r_{i}\cap r_{i+1}$ and no other $r_{j}$ while $x_{k-1}$ belongs to $r_{k-1}\cap r_{0}$ and possibly other hyperedges;
  • Berge-acyclic if it does not contain any Berge cycle. A Berge cycle is a sequence of length $k\geq 2$ of distinct vertices and distinct hyperedges $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ such that $\forall\ i\in[k]:x_{i}\in r_{i}\cap r_{(i+1)\bmod k}$.
The four notions of acyclicity form a strict hierarchy that Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic [13].
Given a hypergraph $H$ with $n$ hyperedges, Algorithm˜1 labels each hyperedge with an integer $i\in[1,n]$ in ascending order and assigns a parent $p(r_{i})$ to each hyperedge $r_{i}$. Upon labeling an edge the algorithm marks all previously unmarked vertices contained in that edge. Initially, all hyperedges are unlabeled and all vertices are unmarked. The algorithm begins by assigning nullnull as the parent of the root $r$ and initializing label $i$ to 0. In each following while loop iteration, the algorithm increments $i$ by one and labels the current hyperedge as $r_{i}$. For each unmarked vertex $x$ in $r_{i}$ (a vertex is considered marked when it is removed from $X$), it assigns $r_{i}$ tentatively as the parent of each unlabeled hyperedge $r^{\prime}$ containing $x$. The algorithm marks all unmarked vertices in $r_{i}$ and removes $r_{i}$ from future consideration. Finally, it selects a hyperedge with the most marked vertices to label next, breaking ties arbitrarily. The algorithm terminates when all hyperedges are labeled. Tarjan and Yannakakis [40] describes an efficient implementation of Algorithm˜1 running in $\mathcal{O}(|H|)$.
Example 42.
Considering the hypergraph $H_{6}$ shown in Figure˜2(a), we run Algorithm˜1 from the hyperedge $P$. It is labeled as $r_{1}$ and all its vertices are marked. $P$ is also tentatively assigned as the parent of five hyperedges such that $p(S)=p(T)=p(U)=p(W)=p(Y)=r_{1}$. We proceed to label the next hyperedge with the most marked vertices. At this point, each unlabeled hyperedge has one marked vertex, so we can choose any of them, say $S$, and label it as $r_{2}$.
Because the labeling of $S$ marks no additional vertices in $T,U,W,Y$, their parents remain unchanged. These remaining hyperedges are again tied with the same number of marked vertices, so we can choose any of them, say $T$, and label it as $r_{3}$. Since the labeling of $T$ marks a vertex $c$ in $U,W$, their parents are updated to $p(U)=p(W)=r_{3}$.
We continue the process until all hyperedges are labeled and all vertices are marked. By connecting each labeled hyperedge with its parent, we construct the join tree $T^{M}$ shown in Figure˜2(d).
Notice that Algorithm˜1 can generate different join trees for the same hypergraph by breaking ties differently. However, it does not guarantee to generate all possible trees. For example, it never generates the Join Tree $T^{G}$ shown in Figure˜2(c).
Appendix B Data Structures for Constant Time Operations on Graphs
Given line graph $L$ and MCS tree $T$, Algorithm˜3 iterates over non-tree edges of $L$. Moreover, the algorithm uses data structures $0pt(\cdot)$, $\mathsf{LA}(\cdot)$ and $\mathsf{LCA}(\cdot)$, in determining the endpoints of a non-tree edge’s LCA edges. In this section we describe briefly how iterating over non-tree edges can be achieved in time linear in the number of vertices and edges of $L$ and how determining the endpoints of a non-tree edge’s LCA edges can be achieved in constant time.
Given a graph $G=(V,E)$, we assume the vertices are represented by integers from 1 to $|V|$. Given a tree $T$ of $G$ rooted at node $r$, recall that data structure $0pt_{T}(\cdot)$, stores for each node $r_{i}$ the length of the shortest path between $r$ and $r_{i}$ in $T$. Data structure $\mathsf{LCA}_{T}(\cdot)$, stores for each pair of nodes $r_{i},r_{j}$ the lowest common ancestor of $r_{i},r_{j}$ in $T$. Finally, data structure $\mathsf{LA}_{T}(\cdot)$ stores for each node-integer pair $r_{i},j$, the ancestor of node $r_{i}$ at depth $j$. Each of these data structures can be established from a tree $T$ in time linear in the size of the tree and support constant time look-ups.
Observation 43.
Given rooted tree $T$, data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$ can be established in time $\mathcal{O}(|T|)$ and support $\mathcal{O}(1)$ time look-ups [9, 33, 3, 4].
Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$ and an MCS tree $T$ of $H$, recall that the LCA edges $\lambda(e)$ of a non-tree edge $e=(r_{i},r_{j})$ of $T$ is defined as the set of edges incident to $\mathsf{LCA}(r_{i},r_{j})$ that lie on the unique path between $r_{i}$ and $r_{j}$ in $T$. The following observation shows that an edges LCA edges can be found in constant time given data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$.
Observation 44.
Given the line graph $L$ of an $\alpha$-acyclic hypergraph $H$, an MCS tree $T$ and data structures $0pt(\cdot)$, $\mathsf{LCA}(\cdot)$ and $\mathsf{LA}(\cdot)$, the LCA edges of any non-tree edge can be found in constant time.
Proof.
Given a non-tree edge $e=(r_{i},r_{j})$ we find $\mathsf{LCA}(r_{i},r_{j})=l$ in constant time by Appendix˜B. The depth $d=0pt(l)+1$ is likewise found in constant time. The edge $e_{1}=(l,r_{1})$, where node $r_{1}=\mathsf{LA}(r_{i},d)$ is found by a constant time look-up, is adjacent to $l$ and lies on the unique path between $r_{i}$ and $r_{j}$. The same holds for edge $e_{2}=(l,r_{2})$ where node $r_{1}=\mathsf{LA}(r_{j},d)$. ∎
In what remains of this section we describe how Algorithm˜3 iterates over and performs sliding transformations on non-tree edges.
We use an array $L$ to represent the adjacency list of a line graph where entry $L[i]$ consists of a linked list of tuples $(j,\omega)$ where $j$ is a neighbor of $i$ and $\omega$ is the weight of the edge $(i,j)$. A rooted tree $T_{r}$ is likewise implemented as an array, where the index $i$ represents the vertex $i$, and the $i$-th entry $T_{r}[i]$ stores $(p,\omega)$ where $p$ is the parent of $i$ and $\omega$ is the weight of the edge $(i,p)$. The depth table of the tree is also implemented as an array to enable constant time lookup for the depth of a vertex $0pt(T_{r},i)$.
Appendix˜C shows how to construct the weighted line graph and the hypergraph given the original query. Then, using the hypergraph we can construct $T_{r}$ and $0pt$ during the MCS algorithm in linear time [40].
As we iterate over each edge $(j,\omega)$ of $L[i]$, we check if the current edge $(i,j)$ is a tree edge by performing constant-time query of $T_{r}[i]$ and $T_{r}[j]$ to see if one of them is the parent of the other. From Appendix˜B we know that we can find the LCA edges $\lambda(i,j)=\{e_{1},e_{2}\}$ in constant time. There’s no general solution to lookup an entry in a linked list in constant time. Instead of iterating over the linked-list representation of $L$ to find the edge weights of $e_{1}$ and $e_{2}$, we can lookup their weights from the array representation of $T_{r}$ in constant time. Let $e_{1}=(r_{1},p(r_{1}))$ and $e_{2}=(r_{2},p(r_{2}))$ with weights $\omega_{1}$ and $\omega_{2}$. Because both $e_{1}$ and $e_{2}$ are tree edges, we can find their weights by querying $T_{r}[r_{1}]=(p(r_{1}),\omega_{1})$ and $T_{r}[r_{2}]=(p(r_{2}),\omega_{2})$. Each of these operations takes constant time. When we need to delete an edge $e^{*}=(i,j)$ from $L$, we first check if $e^{*}$ is a tree edge by checking $T_{r}[i]$ and $T_{r}[j]$. Deleting $e^{*}$ from the linked list takes constant time. Sliding an edge involves updating the endpoints of the edge which takes constant time.
Appendix C From Queries to (Hyper-)Graphs
Algorithm˜3 assumes the weighted line graph and an initial MCS tree are given as input. In practice, we need to construct these graphs from the original SQL query.
We model a query involving $m$ relations as a set of join predicates each in the form of $R_{a}.x_{i}=R_{b}.x_{j}$ such that $a\neq b\wedge a,b\in[m]$, namely no self-joins or duplicates, and assume each relation is uniquely identified by an integer in $[m]$. For each relation $R_{a}$ with arity $k_{a}$, each attribute $x_{i\in[k_{a}]}$ is uniquely identified by an integer in $k_{a}]$. Therefore each predicate is uniquely identified by a tuple $(a,i,b,j)$. Each query $Q$ is represented by a set of such tuples. We will assume that query $Q$ is transatively closed over join predicates, i.e. if $R_{a}.x_{i}=R_{b}.x_{j}$ and $R_{b}.x_{j}=R_{c}.x_{k}$ are in $Q$, then $R_{a}.x_{i}=R_{c}.x_{k}$ is also in $Q$. The tuple $(a,i,b,j)$ is present in $Q$ if and only if $R_{a}.x_{i}=R_{b}.x_{j}$ is a join predicate over relations $R_{a},R_{b}$. We define the size of a query $|Q|$ as the number of join predicates in $Q$, equivalent to the number of $(a,i,b,j)$ tuples. Our goal is to construct the query hypergraph and the weighted line graph from this set of tuples.
C.1 Construction of Hypergraph
We construct the hypergraph $H$ by finding connected components of the predicate graph, which is a graph where there is a vertex for each attribute $R_{a}.x_{i}$, and there is an edge between $R_{a}.x_{i}$ and $R_{b}.x_{j}$ for every predicate $R_{a}.x_{i}=R_{b}.x_{j}$. Each connected component of the predicate graph corresponds to an “equivalence class” of attributes. We then assign a unique variable to each connected component, and map each attribute to the variable of its component. This can be done in linear time using standard algorithms for computing connected components [9, 39].
Algorithm˜4 generates hypergraph $H$ from query $Q$. $H$ is represented as an array of size $m$ where each entry is a list of integers. The algorithm starts by initializing $H$ to an array of size $m$ and $P$ to an array of size $p$, where $p$ is the number of distinct relation-attribute pairs $R_{a}.x_{i}$ in $Q$ as shown by Lines 4-4. If numbers $m$ and $p$ are not known a priori they can be found in a constant number of passes over elements in $Q$. Array $H$ will serve as a hyperedge list of the query hypergraph and array $P$ will serve as an adjacency list representation of the query predicate graph. The subsequent for-loop fills out adjacency list $P$ at Lines 4-4. We assume here that each $(a,i)$ pair is mapped to a unique integer in $[p]$ by a perfect hash function $f$ at Line 4, otherwise such a mapping can be established in a constant number of passes over $Q$. Next, a standard algorithm [9, 39] for the enumeration of connected components of a graph is called in Line 4. This produces array $\mathcal{C}$ where each entry is a list of vertices that forms a connected component in $P$. The subsequent nested for-loop iterates over each relation in each component to fill out hyperedge list $H$ at Lines 4-4.
Input: Query $Q$
Output: Query hypergraph $H$
1
2for $i\in[1\ldots m]$ do
3 $H[i]\leftarrow\texttt{linkedList}()$
4
5for $i\in[1\ldots p]$ do
6 $P[i]\leftarrow\texttt{linkedList}()$
7
8for $(a,i,b,j)\in Q$ do
9 $p=f(a,i)$
10 $P[p].\text{append}(b,j)$
[…]
12
13for $i\in[|\mathcal{C}|]$ do
14 for $R_{a}\in\mathcal{C}[i]$ do
15    $H[a].\text{append}(i)$
16
17
return $H$
Algorithm 4 Generating query hypergraph $H$ from query $Q$.
Observation 45.
Algorithm˜4 generates the hypergraph of $Q$ in time $\mathcal{O}(|Q|)$.
Proof.
Arrays $H$ and $P$ can be initialized in time $\mathcal{O}(|Q|)$ since $m\leq|Q|$ and $p\leq|Q|$. The subsequent for-loop iterates over all elements in $|Q|$ making a single insertion to $P$ thus performing at most $\mathcal{O}(|Q|)$ steps. Predicate graph $P$ therefore has size at most $\mathcal{O}(|Q|)$. Generating a list of connected components of predicate graph $P$ is done in time linear in $|P|=\mathcal{O}(|Q|)$ by standard algorithms [9, 39]. The subsequent nested for-loop iterates over each relation in each component performing a single insertion on each iteration. The sum of the sizes of each component is exactly number of vertices of $P$ and therefore $\mathcal{O}(|Q|)$ elementary operations are performed. ∎
C.2 Construction of Weighted Line Graph
To construct the weighted line graph, we will ignore the attribute indices $(i,j)$ and focus on the relation indices $(a,b)$ for each join predicate in $Q$. The set of predicates can then be seen as representing the edge list of a multigraph. Algorithm˜5 generates the weighted line graph of $Q$ by converting this multigraph into a weighted graph $L$ where the weight of each edge $(u,v)$ is the number of parallel edges between $u$ and $v$ in the multigraph.
The algorithm first initializes arrays $L$ and $M$ both of size $m$, where $M$ will be an adjacency list representation of a multigraph and $L$ will be adjacency list representation of the final weighted line graph in Line 5-5. Auxiliary array weight is initialized in Line 5. Entries of weight are initialized to be an empty linked list. In the subsequent for-loop, "edges" in $Q$ are inserted into multigraph $M$ in Line 5. Array weight[j] stores the weight of the edge $(i,j)$ for the current $i$. Entries of weight[j] are initialized to 0. The subsequent for-loop iterates over adjacency list $M$ in Line 5. The first nested for-loop at Line 5 iterates over each neighbor $j$ of $i$ stored in $M[i]$. For each neighbor $j$, array weight[j] is incremented by 1. Array weight[j] now contains the weights of all edges adjacent to $i$ at its non-zero entries. The second nested for-loop from Line 5 to Line 5 iterates over each neighbor $j$ one more time and inserts neighbor and weight pairs into array $L$ if weight[j] is not 0, otherwise skipping to the next neighbor. After an insertion is made, weight[j] is set to 0. A weight of 0 at weight[j] indicates the edge at $(i,j)$ and its weight has already been outputted. This prevents the insertion of duplicate edges and resets array weight[j] to be reused for the next relation in $M$. Since each array operation takes constant time and we visit each edge a constant number of times, the overall complexity of this procedure is linear in the size of the input.
Input: Query $Q$
Output: Weighted line graph $L$
1
2for $i\in[1\ldots m]$ do
3 $L[i]\leftarrow\texttt{linkedList}()$
4 $M[i]\leftarrow\texttt{linkedList}()$
5 $\texttt{weight}[i]\leftarrow 0$
6
7for $(a,i,b,j)\in Q$ do
8 $M[a].\text{append}(b)$
9
10for $R_{i}\in M$ do
11 for $R_{j}\in M[i]$ do
12    $\texttt{weight}[j]\leftarrow\texttt{weight}[j]+1$
13 for $R_{j}\in M[i]$ do
14    if $\texttt{weight}[j]\neq 0$ then
15       $L[i].\text{append}(R_{j}.\texttt{weight}[j])$
16       $\texttt{weight}[j]\leftarrow 0$
17
18
19
return $L$
Algorithm 5 Generating weighted line graph $L$ from query $Q$.
Observation 46.
Algorithm˜5 generates the weighted line graph of $Q$ in time $\mathcal{O}(|Q|)$.
Proof.
Array initializations for $L$ and $M$ take $\mathcal{O}(m)$ time where $m\leq|Q|$. The first for–loop scans all predicates in $Q$ and inserts the corresponding edges in the multigraph $M$, performing $\mathcal{O}(|Q|)$ insertions, so $|M|=\mathcal{O}(|Q|)$. The auxiliary array weight has size $m$ and is initialized in $\mathcal{O}(m)$ time.
In the subsequent traversal of $M$, each adjacency list is visited once. For every neighbor $j$ of a vertex $i$, the first inner loop performs a constant–time update of weight[j]; the second inner loop performs at most one constant–time insertion of $(j,\texttt{weight[j]})$ into $L$ and resets weight[j]. Thus each stored neighbor (i.e., each edge occurrence in $M$) incurs only $\mathcal{O}(1)$ work overall. Since the total number of stored neighbors is $\Theta(|M|)=\mathcal{O}(|Q|)$, the time for these loops is $\mathcal{O}(|Q|)$.
Summing up, all steps are linear in $|Q|$, hence Algorithm˜5 runs in time $\mathcal{O}(|Q|)$. ∎
Appendix D Missing Proofs in Section˜4
Section˜4.2 If $H^{\prime}\twoheadrightarrow H$ then $\mathcal{T}(H)\subseteq\mathcal{T}(H^{\prime})$.
Proof.
Let $T$ be a join tree of $H$, and $x^{\prime}$ be any vertex in $H^{\prime}$. Consider the neighborhood $H^{\prime}|_{x^{\prime}}$ of $x^{\prime}$ in $H^{\prime}$ comprising the set of hyperedges $\left\{r^{\prime}\in H^{\prime}\mid x^{\prime}\in r^{\prime}\right\}$. Since $H^{\prime}\twoheadrightarrow H$, we have that $x^{\prime}\in r^{\prime}$ if and only if $\mathsf{f}(x^{\prime})\in\mathsf{f}(r^{\prime})=r^{\prime}$. The neighborhood $H|_{\mathsf{f}(x^{\prime})}$ of $\mathsf{f}(x^{\prime})$ in $H$ comprises of the set of hyperedges $\left\{r\in H\mid\mathsf{f}(x^{\prime})\in r\right\}$. Therefore, $r^{\prime}\in H^{\prime}|_{x^{\prime}}\iff\mathsf{f}(r^{\prime})=r^{\prime}\in H|_{\mathsf{f}(x^{\prime})}$. Because $T$ is a join tree of $H$, $H|_{\mathsf{f}(x^{\prime})}$ is connected in $T$, therefore $H^{\prime}|_{x^{\prime}}$ is also connected in $T$ and $T$ is also a join tree of $H^{\prime}$. ∎
Section˜4.2 Let $T$ be an MCS tree. For any edge $e\in T$ that has a parent, then $e\not\subseteq p(e)$.
Proof.
In Algorithm˜1, a child $r_{c}$ is only connected to a parent $r$ if they share some previously unmarked vertex $x$. Because every vertex is marked only once, and the parent $r_{p}$ of $r$ was labeled before $r$, $x$ cannot be marked by $r$’s parent $r_{p}$. Let the $e$ be edge between $r$ and $r_{c}$, then the edge between $r$ and $r_{p}$ is $p(e)$, and $x\in e$ but $x\notin p(e)$, therefore $e\not\subseteq p(e)$. ∎
Section˜4.2 Let $T$ be an MCS tree. For two edges $e,e^{\prime}\in T$ that have parents,
$(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$
Proof.
We prove by contrapositive, assuming $p(e)\neq p(e^{\prime})$.
Case 1:
$p(e^{\prime})=e$, then $(e\setminus p(e))\cap(e^{\prime}\setminus e)\subseteq(e\cap e^{\prime})\setminus e=\emptyset$.
Case 2:
$p(e^{\prime})\neq e$. Without loss of generality, we suppose $e$ is no further from the root than $e^{\prime}$. Let $p(e^{\prime})=e^{\prime\prime}$, then $e^{\prime\prime}$ is on the path between $e$ and $e^{\prime}$ in $T$. If $(e\setminus p(e))\cap(e^{\prime}\setminus e^{\prime\prime})\neq\emptyset$, then $\exists\ x\in(e^{\prime}\cap e)\setminus e^{\prime\prime}$, which violates the running intersection property of join trees.
Section˜4.2 Let $H^{\prime}$ be a duplicated hypergraph of $H$. Then,
  1. 1.
    $H^{\prime}\twoheadrightarrow H$ and $H^{\prime}\twoheadleftarrow H$
  2. 2.
    $\mathcal{T}(H^{\prime})=\mathcal{T}(H)$
  3. 3.
    $L(H^{\prime})=L(H)$
Proof.
1. We only need to show that $H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) holds for $H^{\prime}$ derived from $H$ by a single vertex duplication. $H^{\prime}\twoheadrightarrow H$ ($H^{\prime}\twoheadleftarrow H$) will then hold for any duplicate $H^{\prime}$ derived by a series of vertex duplications since strong homomorphisms are closed under composition.
Suppose $H^{\prime}$ is derived from $H$ by duplicating $x\in X(H)$ to $x^{\prime}$.
Consider a pair of functions $(\mathsf{f}_{X}:X(H^{\prime})\mapsto X(H),R(H)_{\texttt{id}})$ where $\mathsf{f}_{X}$ maps $x^{\prime}$ to $x$. Otherwise, it is an identity map $X(H)_{\texttt{id}}$. This function pair constitutes a strong hypergraph homomorphism $H^{\prime}\twoheadrightarrow H$. Likewise, the pair of functions $(X(H)_{\texttt{id}},R(H)_{\texttt{id}})$ constitutes a homomorphism $H\twoheadrightarrow H^{\prime}$.
2. It follows from Section˜4.2 and the facts above.
3. By the strong homomorphisms, the hyperedge sets have the same cardinality $|R(H)|=|R(H^{\prime})|$. So do the vertex sets of the line graphs, $R(L(H))=R(H)$ and $R(L(H^{\prime}))=R(H^{\prime})$. By construction, the vertex duplication adds $x^{\prime}$ to each hyperedge containing $x$, therefore does not create any additional edges in the line graph, $E(L(H^{\prime}))=E(L(H))$. ∎
Appendix E Missing Proofs in Section˜5
Section˜5 An $\alpha$-acyclic hypergraph is Berge-acyclic if and only if it is linear.
Proof.
Fagin [13] showed that every Berge-acyclic hypergraph is linear. We prove the other direction by contradiction, by assuming that $H$ is $\alpha$-acyclic and linear. We also assume that there is a Berge cycle $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ of length $k\geq 2$. Without loss of generality, we suppose $r_{1}$ is the first hyperedge on the cycle to be removed by GYO reduction. Then $r_{1}$’s parent $r_{p}$ must satisfy $r_{p}\cap r_{1}\supseteq(r_{1}\cap\bigcup_{i\in[k]\wedge i\neq 1}r_{i})=\{x_{0},x_{1}\}$, contradicting $H$’s linearity. ∎
Section˜5.1 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
Proof.
A Berge-acyclic hypergraph is $\gamma$-acyclic [13]. Zhu proves that the line graph of a $\gamma$-acyclic hypergraph is chordal [48]. We now prove that the line graph is diamond-free by showing that every diamond must be part of a $K_{4}$. Assuming a diamond $(r_{1},r_{2},r_{3},r_{4})$ in $L$ without the edge $(r_{1},r_{3})$, it can be treated as two triangles $(r_{1},r_{2},r_{4})$ and $(r_{2},r_{4},r_{3})$ sharing an edge $(r_{2},r_{4})$. We now consider the possible edge labelings in each triangle:
Case 1:
If all edges are labeled with distinct variables, each triangle forms a Berge-cycle, contradicting the Berge-acyclicity.
Case 2:
If two edges are labeled with the same variable, all three vertices share that variable. The third edge must be labeled with the same variable.
Since the two triangles share the edge $(r_{2},r_{4})$, all their edges are labeled with the same variable $x$, namely $x\in r_{1}\cap r_{2}\cap r_{3}\cap r_{4}$. There has to be an edge $(r_{1},r_{3})\in L$, contradicting the assumption that $(r_{1},r_{2},r_{3},r_{4})$ is a diamond. ∎
Section˜5.2 Let $T_{r}$ be the canonical join tree rooted at $r$ of a Berge-acyclic hypergraph. Along its root-to-leaf path $r,r_{1},\ldots,r_{k}$, each pair of adjacent vertices shares a distinct variable.
Proof.
Let $r_{0}=r$, Section˜5 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$. By Theorem˜33, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$. The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is equivalent to the edge label $\chi(r_{i-1},r_{i})$. The edge labels along the path are all distinct. Otherwise, two consecutive edges with the same label allows for shortening the path by removing either, contradicting the shortest path property. Two non-consecutive edges with the same label violate the running intersection property. ∎
Appendix F Additional Figures, Tables and Algorithms
Name # Queries # $\alpha$-Acyclic # Composite-Key Joins # Berge-Acyclic
TPC-H[41] 22 21 2 19
JOB[24] 113 113 0 113
STATS-CEB[17] 2603 2603 0 2603
CE[8] 3004 1839 0 1839
Spider-NLP[46] 4712 4709 6 4703
Table 1: Acyclic queries in the benchmarks (all $\alpha$-acyclic queries are also $\gamma$-acyclic.)
Input: Left-deep plan $r_{1},r_{2},\dots,r_{n}$
Output: Tree rooted at $r_{1}$
1
[…]
3 for $i\leftarrow 2$ to $n$ do
4   key $\leftarrow r_{i}\cap\bigcup r_{k<i}$
5
6 for $j\leftarrow 1$ to i1i-1 do
7    if $\text{key}\subseteq r_{j}$ then
8       $p(r_{i})\leftarrow r_{j}$
9       break
10
11      end if
12
13   end for
14
15 end for
Algorithm 6 Converting a binary plan to a join tree
Figure 5: Possible cases of how $r_{k+1}$ is connected to $G_{k}$
(a) $H_{3}$
(b) $T_{S}(H_{3})$
(c) $H_{5}$
(d) $T_{W}(H_{5})$
(e) $T^{\prime}_{W}(H_{5})$
Figure 6: $H_{3}$ is a non-Berge hypergraph that admits a unique canonical join tree $T_{S}(H_{3})$ with any relation chosen as root, such as $T_{S}(H_{3})$. $H_{5}$ is a $\gamma$ hypergraph that does not admit unique canonical join trees at any relation chosen as root. For example, $T_{W}(H_{5})$ and $T^{\prime}_{W}(H_{5})$ are MCS trees generated by Algorithm˜1. Neither of them is a canonical join tree rooted at $W$.
Input: Hypergraph $H$, root $r$
Output: weighted join tree $T_{r}$ and unweighted line graph $L$
1
[…]
4while $R\neq\emptyset$ do
5 $i\leftarrow i+1;r_{i}\leftarrow r$
6 for $x\in r_{i}\cap X$ do
7    for distinct $r^{\prime},r^{\prime\prime}\in R:x\in r^{\prime}\cap r^{\prime\prime}$ do
8       $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}E(L)\leftarrow E(L)\cup\{(r^{\prime},r^{\prime\prime})\}$
9       $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\omega(r^{\prime},r_{i})\leftarrow|r^{\prime}\cap r_{i}|$
10       $p(r^{\prime})\leftarrow r_{i}$
11
12
13 $X\leftarrow X\setminus r_{i};R\leftarrow R\setminus\{r_{i}\}$
14 $r\leftarrow\operatorname*{arg\,max}_{r\in R}|r\setminus X|$
15
return $L,p(\cdot),\omega(\cdot)$
Algorithm 7 MCS for the $\gamma$-acyclic
Input: weighted MCS tree $T_{r}(H)$ weight function $\omega$ and unweighted union join graph $L(H)$
1 $G^{\equiv}\leftarrow L(H)$
2for $e^{*}=(r_{i},r_{j})\in R(L(H))\setminus R(T_{r}(H))$ do
 // where $0pt(T_{r}(H),r_{i})\leq 0pt(T_{r}(H),r_{j})$
3
4 $l\leftarrow\mathsf{LCA}(r_{i},r_{j});d\leftarrow 0pt(l)+1$
5 $e_{1}=(l,r_{1}\leftarrow\mathsf{LA}(r_{i},d));w_{1}=\omega(e_{1})$
6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
7 if $e_{1}=e_{2}$ then
8      slide $r_{j}$ to $r_{1}$
9
10 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{1}=w_{2}$ then
11      slide $r_{i}$ to $r_{1}$ and $r_{j}$ to $r_{2}$
12
13 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{1}<w_{2}$ then
14      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{1}$
15
16 else if $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}w_{2}<w_{1}$ then
17      slide $r_{i}$ to $l$ and $r_{j}$ to $r_{2}$
18
19return $G^{\equiv}$
Algorithm 8 buildEGbuildEG for the $\gamma$-acyclic